OpenStudy (anonymous):

f g(x) = 2(x − 4), find the value of x if g(x) = 20. 32 12 14 10

OpenStudy (lgbasallote):

g(x) = 2(x-4) g(x) is also = 20 therefore 2(x-4) = 20 distribute 2x - 8 = 20 can you continue from there?

OpenStudy (anonymous):

IS IT 10

OpenStudy (lgbasallote):

no..what did you do to get that?

OpenStudy (anonymous):

14

OpenStudy (lgbasallote):

hmm why don't you try C?